OpenStudy (anonymous):

I'm buying tools for office I have 23 staff and need to get 2 pens and 1 notebook for each of them. The pens are sold in packets of 8 or 10. I decide to buy just the larger packets of each and save any extra ones. I told my assistant how many packets of pens and notebooks to buy but when he returned I found that he had bought the smaller packets in each case. I am going to buy sufficient extra pens and notebooks that I need for my staff singly. A single pen will cost 12 cents and a single notebook will cost 35 cents. How much do I need to pay for the extra pens

OpenStudy (anonymous):

Pens required=23*2=46, Notebooks required=23 Large packets of pen required= 46/10=4.6 packets so I needed to buy 5 large packets and save 4 extra pens. But assistant buys 5 small packets=40 pens. So I still need 6 pens more individually @ 12 cents so cost of 6 extra pens=72 cents Notebooks are neither asked for and nor is there any information given so it's not necessary to calculate it

OpenStudy (anonymous):

\[\text{Ceiling}\left[\frac{23*2}{10}\right]=5 \]\[(23*2)-(5*8)=6 \]\[6*12=72 \text{ cents}\]

OpenStudy (anonymous):

Relient K rules